In the old days a pony express, traveling at the rate of 60 miles a day, had been gone 5 days, when a second pony express was sent after them, travelling at the rate of 75 miles per day. In what time would the one overtake the other?
Answer by fcabanski(1391)

Distance = rate x time


The first rider traveled 60*5 = 300 miles.


The second rider overtakes the first at 75-60 = 15 miles per day and must make up that 300 miles.


300 = 15 * t


300/15 = t = 20 days.

Hope the solution helped. Sometimes you need more than a solution. Contact fcabanski@hotmail.com for online, private tutoring, or personalized problem solving (quick for groups of problems.)